It is planned to build a bridge AB across the river.

Calculate the length of the bridge if the base BC

= 165 m, ß = 108 °, y = 20 °


1
Expert's answer
2022-03-24T18:51:22-0400



To calculate the length of the bridge we use sine rule

"\frac{c}{sineC}=\frac{a}{sineA}"


Therefore,

"\frac{c}{sine20°}=\frac{165°}{sine52°}"


Cross multiply to get the unknown value


"c=\frac{165sine20°}{sine52°}"


"c=152.68m"


Therefore, the length of the bridge will be

"152.68m"